# Compact 3D Scene Representation via Self-Organizing Gaussian Grids

### v1

@@ 4.3 Smoothness Regularization
The magnitude of the smoothness regularization $\lambda$ decides the size of the gradient for local smoothness in the 2D attribute grids in relation to the size of the gradient of the Gaussian training algorithm. To find a good balance between both optimization targets, we tested a series of parameters for $\lambda$. The results for this study is displayed in table 2. As expected, lower magnitudes for $\lambda$ result in higher rendering quality, but also larger storage size.
Neighbor Loss $\lambda$ PSNR Size in MB
@@
To demonstrate the importance of the smoothing regularization, we only apply our sorting method to a normally trained 3DGS model and compress the result. This is demonstrated in table 3. Here, 3DGS + sort denotes the experiment without smoothness regularization. With a size of only 36 MB compared to 174 MB, it is substantially smaller than before. However, the PSNR is also reduced. This is because the compression method relies on quantization and smooth structures, which the sorting algorithm only partially achieves. Both our experiments, on the other hand, both perform better in terms of PSNR and size, underlining the significance of the smoothing regularization.
Method PSNR Size in MB
3DGS 24.89 174.0
3DGS + sort 23.90 36.4
Ours-Q 24.90 17.3
Ours-S 24.63 13.9
Table 3: Evaluating the PSNR and size of a 3DGS model that was not trained with the smoothness regularization. Instead we only apply our sorting method after training.
@@ 4.4 Features to 2D Grid
To convert the millions of Gaussians into a 2D grid, we reshape the features to a square grid. This grid has a side length of the rounded square root of the number of Gaussians. This means that we potentially loose a row and a column of Gaussians at most during the reshape process. To ensure that we do not throw away any Gaussians that might bring meaningful contributions to the scene, we sort all Gaussians by opacity and only discard Gaussians with the lowest opacity. In figure 7, we show that the PSNR of a rendered scene is not affected when removing up to 30% of Gaussians that have a low opacity. In contrast, removing Gaussians with small scaling, affects the quality of the scene considerably more.
@@
4.5 Sorting Performance
To demonstrate the effectiveness of our parallel sorting algorithm, we compare it to the state of the art in quality, FLAS [2]. An implementation of Self-Sorting Maps [33] was not available to us. We measure runtime and, to estimate the quality of the sort, the variance of the absolute differences (VAD) between all neighboring values, over all channels.
Sorting a random 512x512x3 grid on an AMD Ryzen Threadripper PRO 5955WX CPU with FLAS takes 131s and achieves a VAD of 3.53. Our algorithm running on an Nvidia RTX 4090 finishes in 5.7 seconds after 8015 reorders with a VAD of 4.02 (the shuffled data has a VAD of 3607.45).
@@ 5 Conclusion
We have presented a novel optimized representation of 3D Gaussian Splatting, targeting efficient compression and storage. This is achieved by a novel, highly parallelized sorting strategy that smoothly arranges the high dimensional, initially unordered Gaussian parameter sets in a 2D grid, which can be encoded by standard image coding techniques like jpeg-xl. In addition, an optimized training scheme for the 3D Gaussians with a new smoothness loss leads to splat configurations with even higher smoothness in the 2D grid, while still preserving the original accuracy of the 3D scene. In spite of the proposed extensions, training time only negligibly changes compared to the original 3DGS approach. With the proposed solution, we can reduce the data down to 4% of the original size at the same visual quality, enabling the efficient handling of large 3DGS scenes in practical applications.
In future work, we target a further increase in compression efficiency. Besides an adapted weighting of the individual parameters of the smoothness loss, better and less correlated representations for rotations, shape and spherical harmonics will be investigated. The main focus, however, will be on the extension to 4D scenes with consideration of temporal dependencies, supporting the strength of 3D Gaussians in the representation of dynamic scenes.
@@ Acknowledgements
Supplementary Material
6 Results for SSIM
In addition to the results from table 1, we also measure the structural similarity index measure (SSIM) in table 4. Those results correlate with the PSNR results from table 1.
@@ 6 Results for SSIM
3DGS 0.814 607.3 0.866 446.6 0.907 452.8 0.970 55.4
3DGS w/o SH 0.803 54.3 0.859 39.5 0.908 39.9 0.962 5.1
Ours-Q 0.772 23.9 0.838 11.9 0.891 8.4 0.955 2.2
Ours-S 0.763 18.2 0.831 9.9 0.892 7.2 0.953 1.7
Table 4: A comparison over structural similarity index measure (SSIM) between our method, the default 3DGS model and prior NeRF-based methods. Results with a $\dagger$ are directly copied from [14, 20]
7 Parameter Selection
The 3D Gaussian Splatting (3DGS) algorithm comes with many different training parameters. In the following we will give a brief overview of them and highlight the parameters that we changed for our new 3DGS training algorithm.
3DGS Parameters: To minimize the sorting time during training, we make slight changes to the default parameters of 3DGS. I.e. we reduce the number of Gaussians that are created during optimization. To achieve this, we modify the following five parameters:
  • Densification interval $\in\mathbb{N}$: Determines how often the densification process is taking place. A low value corresponds to frequent densification, which results in a large amount of gaussians.
  • Densify grad threshold $\in\mathbb{R}$: All Gaussians with a larger accumulated gradient for the xyz position, are split and cloned. A low threshold results in more Gaussians generated during training.
  • Densify min opacity $\in\mathbb{R}$: A threshold that filters all Gaussians with smaller opacity during densification. A high value leads to fewer Gaussians.
  • Opacity reset interval $\in\mathbb{N}$: In every $x$ steps, the opacity of all Gaussians is set to 0.01.
  • Percent dense $\in\mathbb{R}$: During the densification process, large Gaussians are split into two smaller copies with each 80% of the original size and small Gaussians are cloned identically. The percent dense parameter specifies the threshold at which a gaussian is classified as small or large. The value, which set is between 0 and 1, is multiplied by the extend of the scene and used as comparison.
Smoothness Reg. Parameters: The smoothness regularization first blurs the 2D attribute grids and then computes a error term with a Huber loss. This gives us several new parameters to tune:
@@ 7 Parameter Selection
  • Overall multiplier $\lambda\in\mathbb{R}$: A multiplier that scales the loss of the Smoothness Regularization before adding it to the loss of 3DGS.
  • Separate multiplier $\in\mathbb{R}$: A loss scalar for each of the five (position, color, opacity, scaling and rotation) 2D attribute grids.
Sorting Parameters: The sorting algorithm that takes place after every densification can be adjusted with separate multipliers for each of the five 3DGS attributes (position, color, opacity, scaling and rotation).
In table 5 we give a detailed description of our parameter selection for our two different configurations.
Parameter / Method 3DGS Ours-S Ours-Q
3DGS
Densification interval 100 1000 1000
Densify grad threshold 0.0002 0.00009 0.00007
Densify min opacity 0.005 0.1 0.1
Opacity reset interval 3000 $\infty$ $\infty$
Percent dense 0.01 1.0 0.1
Smoothness Reg.
Kernel size - 5 5
Sigma - 4 3
Overall multiplier $\lambda$ 1.0 1.0
Position weight - 0.0 0.0
Color weight - 0.0 0.0
Opacity weight - 0.33 0.09
Scaling weight - 0.0 0.0
Rotation weight - 0.67 0.91
Sorting
Position weight - 1.0 1.0
Color weight - 1.0 1.0
Opacity weight - 0.0 0.0
Scaling weight - 1.0 1.0
Rotation weight - 0.0 0.0
8 Sorting performance
@@
Figure 8: Behavior of the sorting algorithm under different break thresholds. Decreasing the relative L2 threshold trades off additional sorting quality for longer runtime. These values were measured with an NVidia RTX 4090 on a random 512x512x3 grid. In all our training experiments, we have fixed this parameter to $10^{-4}$.
The parameter that has the largest influence on the grid sorting performance is the threshold, which decides whether the relative reduction in L2 distance between two iterations should stop sorting with the current configuration, and potentially continue with the next permutation or next lower level. Decreasing it leads to a higher sorting quality, but the increase in iterations takes additional runtime. In Figure 8 we have plotted the sorting quality over the total runtime.
@@
In Figure 9, we are showing the runtime performance of the sorting algorithm depending on the input grid size. With the chosen parameters, sorting will usually take less than 10 seconds during training. The highest number of Gaussians of any of our models over all datasets is the Garden scene with 4.37M Gaussians, which requires a grid of a side length of 2091. Thus even the largest of scenes in the used datasets can be sorted in well below a minute.
9 Quantization & Compression
Once training has finished, we can use off-the-shelf image compression methods to store the data on disk.
We store the un-activated Gaussian parameters, retaining the activation methods from 3DGS: sigmoid actiation for opacity and exponential activation for the scale. As mentioned in Section 3, we introduce an additional exponential activation for the coordinates $x$ used during training, allowing for more relative precision in coordinates in the center of the scene:
$x=\text{sign}(x_{\text{log}})\times(\exp(|x_{\text{log}}|)-1)$ (2)
We clip the RGB values (DC values of the spherical harmonics) to the range $[-2,4]$ and normalize it to the range $[0,1]$. Opacity values are clipped to the range $[-6,12]$ and similarly normalized. These ranges are chosen to cover the 1st to 99th percentile of these values across all models from the 360 dataset.
Name PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Size (MB) $\downarrow$
PLY 25.41 0.880 0.152 104.89
NPZ 25.41 0.880 0.152 76.10
JXL ll 25.41 0.880 0.152 66.05
PNG 16 24.08 0.873 0.157 37.00
EXR 25.36 0.878 0.154 30.56
EXR+JXL q 24.89 0.862 0.170 17.25
We quantize the different attributes by rounding them to the closest value of a linear range of $q$ total values, with $q_{\text{coords}}=2^{13}$, $q_{\text{scale}}=q_{\text{rotation}}=2^{6}$ and $q_{\text{opacity}}=2^{5}$. Clipping, normalization and quantization are done post-training in our method. In future work, it would be interesting to investigate applying them during training for potentially further improving the results.
We compress the RGB grid with JPEG XL, as an 8-Bit image with a quality level of 90, and store all other attributes as 32-Bit OpenEXR images with zip compression.
Table 6 contrasts different compression methods on the Truck scene. Described in the text above was the EXR+JXL q method, which we used as the default method for all data sets in the comparisons in the paper.
Notably, the PLY file when trained with our method is of smaller size and of higher quality than training with the default 3DGS parameters (which yields 24.90 PSNR at 174 MB). NPZ is using NumPy’s save_compressed method to directly store the tensors. JXL ll stores all tensors as lossless JPEG XL images. PNG 16 truncates the value below the 1st and above the 99th percentile (over all 360 datasets), then normalizes the data to the $[0,1]$ range and stores them as 32-Bit PNGs. EXR stores the values as 32-Bit OpenEXR files with zip compression. Even though we are not using spherical harmonics, this is providing close to the same quality as 3DGS with spherical harmonics (25.44 PSNR at 615 MB), at a compression rate of over 20x.
In our work, we have demonstrated the feasibility of organizing Gaussians into 2D grids for a compact representation. We have not used spherical harmonics (SH), as they have a relatively small effect on scene quality, while taking three quarters of all space in vanilla 3DGS (45 of 59 attributes in 3DGS are spherical harmonics, with the default third degree SH). We believe that a better space-for-quality trade-off is required for efficient representation of view-dependent effects. When this is identified in future work, we think it will be worthwhile to perform additional fine-tuning of the quantization and compression parameters.

### v2

@@ 4.1 Smoothness Regularization
The magnitude of the smoothness regularization $\lambda$ decides the size of the gradient for local smoothness in the 2D attribute grids in relation to the size of the gradient of the Gaussian training algorithm. To find a good balance between both optimization targets, we tested a series of parameters for $\lambda$, see table 3. As expected, lower magnitudes for $\lambda$ result in higher rendering quality, but also larger storage size.
$\lambda$ PSNR Size in MB
@@
w/ SH w/o SH
Method PSNR Size PSNR Size
Ours 25.37 34.3 24.90 17.3
3DGS Our Compression 21.98 71.7 20.37 44.5
3DGS .ply 25.44 624.7 24.89 174.0
3DGS .ply.zip 25.44 548.2 24.89 130.8
Blog Post [28] 25.05 41.5 - -
Table 3: Comparison of our method to other 3DGS compression approaches, calculated using the Truck scene from Tanks&Temples. Size is in MB.
4.2 Synergy of Sorting and Compression
To validate the effectiveness of the sorting and smoothing during the training, we measure the storage size and PSNR of a model that is only compressed after the vanilla 3DGS training. The result is shown in table 3 3DGS Our Compression. We observe a much lower PSNR compared to our result with sorting and smoothing activated. This underlines that the local smoothness, caused by sorting and smoothing, plays a significant role for the success of our approach.
To put our approach into relation with available compression methods we perform two further comparisons in Table 3. Firstly, we simply compress the 3DGS .ply file with a zip compression. And secondly, we compare the results of the method developed in the blog post series Making Gaussian Splats smaller [28]. As a result, we observe that our method outperforms the .zip compression substantially, reducing the storage size by a factor of 16x. For the comparison with Making Gaussian Splats smaller, the results are closer, nevertheless, our method achieves both higher PSNR and lower storage size.
It may be possible to encode the Gaussian Splats with conventional point cloud compression methods. The popular DRACO algorithm [10] supports encoding additional attributes on top of position and color in theory, but its implementations do not [22]. Thus, a direct comparison is unfortunately not possible. Our method’s strength lies in adapting the representation during training to be well compressible (see Section 4.2). Using any post-training compression method, like DRACO, cannot make use of these gains.
Name PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Size (MB)
PLY 25.41 0.880 0.152 104.89
NPZ 25.41 0.880 0.152 76.10
JPEG XL lossless 25.41 0.880 0.152 66.05
PNG 16 24.08 0.873 0.157 37.00
EXR 25.36 0.878 0.154 30.56
JPEG XL quantized 25.14 0.870 0.161 11.86
4.3 Choice of compression format
Table 4 contrasts different compression methods on the Truck scene, without using spherical harmonics. The scene was trained with our parameters and sorted into 2D grids. Then we sampled different image and array encoding formats, measuring rendering quality and file size.Of the lossless formats, JPEG XL provides the smallest files. Notably, the PLY file when trained with our method is of smaller size and of higher quality than training with the default 3DGS parameters (which yields 24.90 PSNR at 174 MB). NPZ is using NumPy’s save_compressed method to directly store the tensors. JXL ll stores all tensors as lossless JPEG XL images.
By introducing lossy compression, the file size can be further decreased, with different tradeoffs between rendering quality and storage space used. PNG 16 truncates the value below the 1st and above the 99th percentile (over all 360 datasets), then normalizes the data to the $[0,1]$ range and stores them as 32-Bit PNGs. EXR stores the values as 32-Bit OpenEXR files with zip compression. We have found using JPEG XL with different levels of quantization provides the best tradeoff. Even though we are not using spherical harmonics, our chosen compression method is providing close to the same quality as 3DGS with spherical harmonics (25.44 PSNR at 615 MB), at a compression rate of 52x.
@@ 4.4 Features to 2D Grid
To convert the millions of Gaussians into a 2D grid, we reshape the list of features into a square grid. As the number of Gaussians grows over time from the densification process, so will the the size of the grid. For simplicity, we find the largest square grid size that can be completely filled with Gaussians, and discard the Gaussians with the lowest opacity values that will not fit. Losing up to one row and one column of Gaussians does not meaningfully alter the scene: In figure 8, we show that the PSNR of a rendered scene is not affected when removing up to 30% of Gaussians with the lowest opacity. In contrast, removing Gaussians with small scaling, affects the quality of the scene considerably more.
@@ 5 Conclusion
We have presented an optimized representation of 3D Gaussian Splatting, targeting efficient compression and storage. This is achieved by a novel, highly parallelized sorting strategy that smoothly arranges the high dimensional, initially unordered Gaussian parameter sets in a 2D grid, which can be encoded by standard image coding techniques like JPEG XL. In addition, an optimized training scheme for the 3D Gaussians with a new smoothness loss leads to splat configurations with even higher smoothness in the 2D grid, while still preserving the original accuracy of the 3D scene. In spite of the proposed extensions, training time only negligibly changes compared to the original 3DGS approach. With the proposed solution, we can reduce the data down to 2% of the original size at the same visual quality, enabling the efficient handling of large 3DGS scenes in practical applications.
In future work, we aim to further enhance compression efficiency. Besides an adapted weighting of the individual parameters of the smoothness loss, better and less correlated representations for rotations, shape and spherical harmonics will be investigated. To achieve even better results, it would be interesting to perform the quantization during training, which is currently happening only once before the compression. The main focus, however, will be on the extension to 4D scenes with consideration of temporal dependencies, supporting the strength of 3D Gaussians in the representation of dynamic scenes.
@@ Acknowledgements
Appendix
Appendix A Results for SSIM
In addition to the results from table 1, we also measure the structural similarity index measure (SSIM) in table 5. These results correlate with the PSNR results from table 1.
@@ Appendix A Results for SSIM
3DGS 0.814 785 0.866 454 0.907 699 0.970 71.6
3DGS w/o SH 0.803 212 0.859 122 0.908 191 0.962 19.4
Ours 0.814 40.3 0.864 21.4 0.909 16.8 0.969 4.1
Ours w/o SH 0.803 16.7 0.857 8.2 0.908 5.5 0.961 2.0
Table 5: A comparison over structural similarity index measure (SSIM) between our method, the default 3DGS model and prior NeRF-based methods. Results with a $\dagger$ are directly copied from [15, 21]. Our method achieves close to the same SSIM as vanilla 3DGS, while reducing size with a factor of 17x to 42x, depending on the dataset.
Appendix B Parameter Selection
The 3D Gaussian Splatting (3DGS) algorithm involves many different training parameters. In the following, we will give a brief overview on these parameters and highlight those we changed for our new training algorithm.
B.1 Densification
To minimize the sorting time during training, we make minimal changes to the default parameters of 3DGS, i.e. we reduce the number of Gaussians that are created during optimization. To achieve this, we modify the following five parameters:
  • Densification interval $\in\mathbb{N}$: Determines how often the densification process executed. A low value corresponds to frequent densification, which results in a large amount of Gaussians.
  • Densify grad threshold $\in\mathbb{R}$: All Gaussians with a larger accumulated gradient for the xyz position are split and cloned. A low threshold results in more Gaussians generated during training.
  • Densify min opacity $\in\mathbb{R}$: This threshold filters all Gaussians with smaller opacity during densification. A high value leads to fewer Gaussians.
  • Opacity reset interval $\in\mathbb{N}$: Every $x$ steps, the opacity of all Gaussians is set to 0.01.
  • Percent dense $\in\mathbb{R}$: During the densification process, large Gaussians are split into two smaller copies, with 80% of the original size each, while small Gaussians are cloned identically. The percent dense parameter specifies the threshold at which a Gaussian is classified as small or large. The value, which is set between 0 and 1, is multiplied by the extend of the scene and used as comparison.
Parameter / Method 3DGS Ours
3DGS
Densification interval 100 1000
Densify grad threshold $2\times 10^{-4}$ $7\times 10^{-5}$
Densify min opacity 0.005 0.1
Opacity reset interval 3000 $\infty$
Percent dense 0.01 0.1
Smoothness Reg.
Kernel size - 5
Sigma - 3
Overall multiplier $\lambda$ 1.0
Weights
Position - 0.0
Color (SH DC) - 0.0
Opacity - 0.09
Scaling - 0.0
Rotation - 0.91
Sph. Harmonics (rest) - 0.0
Sorting weights
Position - 1.0
Color (SH DC) - 1.0
Opacity - 0.0
Scaling - 1.0
Rotation - 0.0
Sph. Harmonics (rest) - 0.0
Table 7: By changing 3DGS default parameters that control the densification step, we observe a different number of Gaussians during the training. Our method creates less Gaussian splats and does not apply regular pruning, which happens every 3k steps with the default parameters.
B.2 Smoothness Regulation
The smoothness regularization first blurs the 2D attribute grids and then computes an error term with a Huber loss. This results in new parameters to tune:
@@ B.2 Smoothness Regulation
  • Overall multiplier $\lambda\in\mathbb{R}$: A multiplier scaling the loss of the Smoothness Regularization before adding it to the loss of 3DGS.
  • Separate multiplier $\in\mathbb{R}$: A loss scalar for each of the five 2D attribute grids (position, color, opacity, scaling and rotation).
B.3 Sorting
The sorting algorithm that takes place after every densification can be adjusted with separate multipliers for each of the 3DGS attributes: position, spherical harmonics (DC and rest), opacity, scaling and rotation. In Table 7, we contrast the vanilla 3DGS parameters with the ones we chose.
B.4 Hyperparameter sensitivity
We found our current sets of parameters through an iterative search, optimizing with the goal to stay close to vanilla 3DGS quality, while minimizing file size. Our guideline for the found parameters is the following: position, color and scale have a large effect on the PSNR, thus they were selected as the sorting keys. Opacity and rotation have a much smaller effect. Therefore, they are smoothed in their neighborhoods instead (with neighbors defined by position/color/scale), to be compressible.
We applied the same parameter sets for the evaluation across all datasets, including synthetic ones, demonstrating their generalizability. Superior configurations might be uncovered through an exhaustive parameter sweep. We think performing this computationally expensive work is best done in future, after choosing a more compact representation than the high-dimensional spherical harmonics attributes, for view-dependent effects.
Appendix C Sorting performance
@@
Figure 9: Behavior of the sorting algorithm under different break thresholds. Decreasing the relative L2 threshold trades off additional sorting quality for longer runtime. These values were measured with an NVidia RTX 4090 on a random 512x512x3 grid. In all our training experiments, we fixed this parameter to $10^{-4}$.
@@
To demonstrate the effectiveness of our parallel sorting algorithm, we compare it to the state of the art in quality, FLAS [2]. An implementation of Self-Sorting Maps [36] was not available to us. We measure runtime and, to estimate the quality of the sort, the variance of the absolute differences (VAD) between all neighboring values, over all channels. Sorting a random 512x512x3 grid on an AMD Ryzen Threadripper PRO 5955WX CPU with FLAS takes 131s and achieves a VAD of 3.53. Our algorithm running on an Nvidia RTX 4090 finishes in 5.7 seconds after 8015 reorders with a VAD of 4.02 (the shuffled data has a VAD of 3607.45). We therefore attain a sorting quality comparable to that of FLAS but with a significantly reduced runtime, thanks to our algorithm’s high degree of parallelism.
The parameter that has the largest influence on the grid sorting performance is the threshold, which decides whether the relative reduction in L2 distance between two iterations should stop sorting with the current configuration, and potentially continue with the next permutation or next lower level. Decreasing it leads to a higher sorting quality, but the increase in iterations takes additional runtime. In Figure 10, we have plotted the sorting quality over the total runtime.
In Figure 10, we show the runtime performance of the sorting algorithm depending on the input grid size. With the chosen parameters, sorting will usually take less than 10 seconds during training. The highest number of Gaussians of any of our models over all datasets is the Garden scene with 4.37M Gaussians, which requires a grid of a side length of 2091. Thus, even the largest of scenes in the used datasets can be sorted in well below a minute.
Table 3 suggests that the visual quality is drastically reduced when only applying sorting and compression after training a vanilla 3D Gaussian Splatting model. This can also be observed visually in Figure 11. This underlines that the smoothing regularization, which influences the 3DGS training, plays a significant role for compressing the attributes efficiently.
Figure 11: A comparison between the rendering quality when only applying our compression method to the vanilla 3DGS model without smoothing.